# Regular Lagrangians in Lefschetz fibrations

### v1

@@ 1.1. Background
In this paper we study regularity in the context of Lefschetz fibrations. As the holomorphic analogue of Morse functions, Lefschetz fibrations bridge complex and symplectic geometry [Don96, Don99, Sei08, MS10], are instrumental in filling classifications in low-dimensional contact topology [Wen10, PVHM10, Öz15], and have generally led to great insight in Weinstein topology [McL09, GP17, CM19].
@@ 1.3. Applications
It is then natural to ask which exact fillable Legendrian links admit such fillings with associated CAL-skeleta. Without prescribing the Legendrian isotopy type, we can answer in complete generality for topological link types using results of Casals [Cas22], Boileau and Orevkov [BO01], and Rudolph [Rud83]:
Corollary 1.1.
Given any exact fillable Legendrian link $\Lambda$, there exists a Legendrian link $\Lambda^{\prime}$, smoothly isotopic to $\Lambda$, such that $\Lambda^{\prime}$ has a regular filling $L$ with an associated CAL-skeleton $L\cup\mathcal{D}$ for $(B^{4},\Lambda^{\prime})$. Further, there is a Lefschetz fibration on $B^{4}$ which is admissible for $L$ and all mutations of $L$ along disks in $\mathcal{D}$.
Since the filling $L$ in the statement above is regular, Corollary 1.1 shows that every exact fillable topological link type in $S^{3}$ is in fact regularly fillable, answering a question in [Bre24]. In higher dimensions, this follows from Lazarev’s existence $h$-principle for regular Lagrangians [Laz20]. In fact, in analogy to Lazarev’s $h$-principle result, one could ask the following question.
@@
For Legendrian links of this form, exact Lagrangian fillings have been constructed using decomposable cobordisms [EHK12, Pan16], Legendrian weaves [TZ18, CZ22, CG24, CGGS24], and conjugate fillings [STWZ19, CL22]. It is shown [CL22, Hug23] that all these constructions are Hamiltonian isotopic. Further, it is known [CW24] that the sheaf moduli $\mathcal{M}_{1}(\Lambda_{\beta})$ admits a cluster structure, and every cluster seed of the cluster structure has a corresponding filling [CG24]. Specifically, Casals and Gao use an initial filling $L_{\mathrm{init}}$ constructed using conjugate fillings, and show that all the other fillings can be obtained by the process of Lagrangian mutations. Conjecturally, all fillings of $\Lambda_{\beta}$ are obtained in this way. We can summarize the above discussion in the following:
@@ 1.4. Organization
In Section 2 we collect background information on coupled Weinstein handles, regular Lagrangians, and open book decompositions. In Section 3 we review and reformulate the work of Johns [Joh11] and S. Lee [Lee25] in a framework we call coupled Lefschetz handles, and in Section 4 we rephrase and prove Theorem 1.2. Finally, in Section 5 we apply the above results to visualizing exact Lagrangian fillings of Legendrian links in $(S^{3},\xi_{\mathrm{st}})$. We build a particular Lefschetz fibration compatible, in the sense of Theorem 1.2, with a filling $L$ and its associated CAL-skeleton. We then show that fillings obtained by Lagrangian mutations are also compatible with the Lefschetz fibration, proving Theorem 1.3. We finish by constructing fillings with associated CAL-skeleta, and proving Corollary 1.1, Proposition 1.1 and Corollary 1.2.
@@ 1.5. Acknowledgments
The authors thank James Hughes and Oleg Lazarev for interest in our work and comments on an early draft. We additionally thank Orsola Capovilla-Searle, Austin Christian, John Etnyre, and Ko Honda for helpful discussions.
@@ 5. Mutations via arc projections
Let $L$ be a properly embedded exact Lagrangian in a symplectic manifold $M$, and $D$ be an embedded Lagrangian disk in the interior of $M$ such that $L\cap D=\partial D$ cleanly, i.e. $D$ is transverse to $L$ along $\partial D$ and $\partial D$ is an embedded smooth curve in $L$. Such a disk is called an $\mathbb{L}$-compressing disk. [^1] This definition was coined by Casals and Weng in [CW24]. Then, the mutated Lagrangian $\mu_{D}(L)$ is obtained by Lagrangian disk surgery (see [Pol91, Yau17]) of $L$ along $D$.
It is shown in [Yau17] that if $L$ is exact, so is $\mu_{D}(L)$, and they are smoothly isotopic. In addition, $\mu_{D}(L)$ has an $\mathbb{L}$-compressing disk $D^{\prime}$ such that $\mu_{D^{\prime}}(\mu_{D}(L))=L$. The argument of [CG24] and a number of related papers in the area (e.g. [CL22, CW24, CGGS24]) use the notion of an $\mathbb{L}$-compressing system, which is a maximal linearly independent (in $H_{1}(L)$) collection of curves that simultaneously bound $\mathbb{L}$-compressing disks.
@@ 5.2. CAL-skeleta associated to fillings
We now recall Casals’ result for Legendrian representatives of algebraic links. For Legendrian links which are the max-tb representatives of the link of an isolated plane curve singularity, [Cas22] provides a construction of a closed arboreal skeleton for $(B^{4},\Lambda)$. Specifically, the main result states that given a Legendrian link $\Lambda$ whose transverse pushoff is the link of an isolated singularity $f\in\mathbb{C}[x,y]$, there exists a Weinstein structure $\mathcal{W}$ on $(B^{4},\omega_{\mathrm{st}})$ such that its skeleton is formed of
  1. (1)
    an exact Lagrangian filling $L$ of $\Lambda$ homeomorphic to a Milnor fiber of $f$, and
  2. (2)
    Lagrangian disks $D_{i}$ attached to $L$ along simple closed curves $\gamma_{i}$ corresponding to vanishing cycles of the Weinstein structure associated to a real Morsification $\tilde{f}$ of $f$.
Theorem 5.1.
[Cas22, Theorem 1.1] Let $f\in\mathbb{C}[x,y]$ be an isolated plane curve singularity and $\Lambda(f)\in(S^{3},\xi_{\mathrm{st}})$ its associated Legendrian link. The Weinstein pair $(B^{4},\Lambda(f))$ admits the arboreal Lagrangian skeleton $\mathbb{L}(f)=\mathfrak{M}_{f}$, obtained by attaching the Lagrangian $D^{2}$-thimbles of $\tilde{f}$ to the Milnor fiber $M_{\tilde{f}}$, for any real Morsification $\tilde{f}\in\mathbb{R}[x,y]$.
Proof of Corollary 1.1.
If $\Lambda$ is exact Lagrangian fillable, its transverse pushoff $T(\Lambda)$ is symplectically fillable [CGHS14, Lemma 4.1]. It follows from [BO01, Theorem 1.2] that $T(\Lambda)$ is quasipositive. The main result of [Rud83] then says that $T(\Lambda)$ is an algebraic link. Consequently, Theorem 5.1 shows that $T(\Lambda)$ has an exact fillable Legendrian representative $\Lambda^{\prime}$ where the filling, denoted $L$, has an associated CAL-skeleton $L\cup\mathcal{D}$. By Theorem 1.2, there is an admissible Lefschetz fibration $p$ on $B^{4}$ for $(B^{4},L)$. Applying Proposition 5.1, any filling of $\Lambda^{\prime}$ obtained by a sequence of mutations along disks in $\mathcal{D}$ is also arc-admissible for $p$. ∎

### v2

@@ 1.1. Background
In this paper we study regularity in the context of Lefschetz fibrations. As the holomorphic analogue of Morse functions, Lefschetz fibrations bridge complex and symplectic geometry [Don96, Don99, Sei08, MS10], are instrumental in filling classifications in low-dimensional contact topology [Wen10, PVHM10, Ozb15], and have generally led to great insight in Weinstein topology [McL09, GP17, CM19].
@@ 1.3. Applications
It is natural to ask which exact fillable Legendrian links admit such fillings with associated CAL-skeleta, or more generally which exact fillable links admit regular fillings (as asked in [Bre24]). The question of whether an exact filling of a Legendrian link is part of a CAL-skeleton also has consequences to associated cluster structures. In this case, work of Casals and Li [CL24] ensures that microlocal merodromies along positive relative cycles on the filling extend to global regular functions on the sheaf moduli of the boundary Legendrian. In particular, this allows for the symplectic construction of cluster structures on these sheaf moduli (i.e., the main results of [CW24a] and [CGG+25]) without relying on combinatorics. In analogy to Lazarev’s $h$-principle result in higher dimensions [Laz20], we formulate the following question.
@@
A weaker question is the following:
Question 1.2.
Given any exact fillable Legendrian link $\Lambda$, does there exists a Legendrian link $\Lambda^{\prime}$, smoothly isotopic to $\Lambda$, such that $\Lambda^{\prime}$ has a regular filling $L$ with an associated CAL-skeleton $L\cup\mathcal{D}$ for $(B^{4},\Lambda^{\prime})$?
Work of Casals [Cas22, Theorem 1.1] establishes this affirmatively for (the unique) max-tb Legendrian representatives of links of single algebraic singularities; such links are iterated cables of the unknot. Combining work of Boileau and Orevkov [BO01] and Rudolph [Rud83], every exact fillable topological link type is algebraic, in that it bounds a smooth algebraic curve in $\mathbb{C}^{2}$. Thus, one approach to Question 1.2 is to extend Casals’ work on links of single singularities to the more general class of algebraic links as described in Rudolph [Rud83].
@@
For Legendrian links of this form, exact Lagrangian fillings have been constructed using decomposable cobordisms [EHK12, Pan16], Legendrian weaves [TZ18, CZ22, CG24, CGGS24], and conjugate fillings [STWZ19, CL22]. It is shown [CL22, Hug23] that all these constructions are Hamiltonian isotopic. Further, it is known [CW24b] that the sheaf moduli $\mathcal{M}_{1}(\Lambda_{\beta})$ admits a cluster structure, and every cluster seed of the cluster structure has a corresponding filling [CG24]. Specifically, Casals and Gao use an initial filling $L_{\mathrm{init}}$ constructed using conjugate fillings, and show that all the other fillings can be obtained by the process of Lagrangian mutations. Conjecturally, all fillings of $\Lambda_{\beta}$ are obtained in this way. We can summarize the above discussion in the following:
@@ 1.4. Organization
In Section 2 we collect background information on coupled Weinstein handles, regular Lagrangians, and open book decompositions. In Section 3 we review and reformulate the work of Johns [Joh11] and S. Lee [Lee25] in a framework we call coupled Lefschetz handles, and in Section 4 we rephrase and prove Theorem 1.2. Finally, in Section 5 we apply the above results to visualizing exact Lagrangian fillings of Legendrian links in $(S^{3},\xi_{\mathrm{st}})$. We build a particular Lefschetz fibration compatible, in the sense of Theorem 1.2, with a filling $L$ and its associated CAL-skeleton. We then show that fillings obtained by Lagrangian mutations are also compatible with the Lefschetz fibration, proving Theorem 1.3. We finish by constructing fillings with associated CAL-skeleta, and proving Proposition 1.1 and Corollary 1.1.
@@ 1.5. Acknowledgments
The authors thank James Hughes and Oleg Lazarev for interest in our work and comments on preliminary draft. We thank Eduardo Fernandez and Kyle Hayden for helpful observations concerning an earlier version of this work related to Question 1.2. We additionally thank Orsola Capovilla-Searle, Austin Christian, John Etnyre, and Ko Honda for helpful discussions.
@@ 5. Mutations via arc projections
Let $L$ be a properly embedded exact Lagrangian in a symplectic manifold $M$, and $D$ be an embedded Lagrangian disk in the interior of $M$ such that $L\cap D=\partial D$ cleanly, i.e. $D$ is transverse to $L$ along $\partial D$ and $\partial D$ is an embedded smooth curve in $L$. Such a disk is called an $\mathbb{L}$-compressing disk. [^1] This definition was coined by Casals and Weng in [CW24b]. Then, the mutated Lagrangian $\mu_{D}(L)$ is obtained by Lagrangian disk surgery (see [Pol91, Yau17]) of $L$ along $D$.
It is shown in [Yau17] that if $L$ is exact, so is $\mu_{D}(L)$, and they are smoothly isotopic. In addition, $\mu_{D}(L)$ has an $\mathbb{L}$-compressing disk $D^{\prime}$ such that $\mu_{D^{\prime}}(\mu_{D}(L))=L$. The argument of [CG24] and a number of related papers in the area (e.g. [CL22, CW24b, CGGS24]) use the notion of an $\mathbb{L}$-compressing system, which is a maximal linearly independent (in $H_{1}(L)$) collection of curves that simultaneously bound $\mathbb{L}$-compressing disks.